Solve the following system of equations using matrices: x+y+z= -9, x-y+5z= -19, 5x+y+z= -29.
Answer by Alan3354(69443)

Solve the following system of equations using matrices:
x+y+z= -9, x-y+5z= -19, 5x+y+z= -29.
---------
|+1 +1 +1 -9|
|+1 -1 +5 -19|
|+5 +1 +1 -29|
----------
Det = 1(-1-5) -1(1-25) +1(1+5) = -6 + 24 + 6
Det = 24
---------
x*Det = 1(-145+19) -1(29+19) -9(-1-5) = -126 - 48 + 54 = - 120
x = -5
------
-y*Det = 1(-145+19) -1(-29+95) -9(1-25) = -126 - 66 + 216 = 24
y = -1
--------
z*Det = 1(29+19) -1(-29+95) -9(1+5) = 48 - 66 - 54 = -72
z = -3